# Fast quantum gates for exchange-only qubits using simultaneous exchange pulses

### v1

@@ Abstract
The benefit of exchange-only qubits compared to other spin qubit types is the universal control using only voltage controlled exchange interactions between neighboring spins. As a compromise, qubit operations have to be constructed from non-orthogonal rotation axes of the Bloch sphere and result in rather long pulsing sequences. This paper aims to develop a faster implementation of single-qubit and two-qubit gates using simultaneous exchange pulses. We introduce pulse sequences in which single-qubit gates could be executed faster and show that subsequences on three spins in two-qubit gates could be implemented in fewer steps. Our findings can particularly speed up gate sequences for realistic idle times between sequential pulses and we show that this advantage increases with more interconnectivity of the quantum dots. We further demonstrate how a phase operation can introduce a relative phase between the computational and some of the leakage states, which can be advantageous for the construction of two-qubit gates. In addition to our theoretical analysis, we experimentally demonstrate and characterize a simultaneous exchange implementation of $X$ rotations in a SiGe quantum dot device and compare to the state of the art with sequential exchange pulses.
@@ I Introduction
In the EO architecture, each qubit is represented by the spins of three electrons hosted in three adjacent quantum dots (Fig. 1(a)), and quantum gates are generated by pulsing exchange interactions between electrons. Out of the resulting eight-dimensional Hilbert space, four states are computational states and four are leakage states. Single-qubit gates involve exchange interactions between the electrons belonging to the same qubit and are usually composed of up to four rotations around two different axes [7, 9] (Fig. 1(b)). Two-qubit gates require exchange interactions between electrons from different qubits. Non-trivial pulse sequences of several exchange pulses populating and depopulating leakage states result in gates such as the CNOT, CZ or SWAP gate [7, 10, 11, 12, 13]. With few exceptions, the sequences that are found to result in two-qubit gates, however, assume only commuting exchange pulses to be pulsed at a time. Here we investigate simultaneous application of non-commuting exchange-pulses to obtain quantum gates with fewer pulses, and hence faster executions of gate sequences. We derive conditions on the exchange signals and compare the results to sequential gate operations.
Figure 1: Schematic of an exchange-only qubit encoded in 3 linearly connected quantum dots in the (1,1,1) charging regime formed by potentials of plunger gates (${\rm P}_{i}$ of dot $i$) and barrier gates (${\rm X}_{ij}$ between dots $i$ and $j$). (a) Qubits are initialized in the singlet state of spins 1 and 2. Single qubit rotations are realized by switching on the exchange interaction (i) $J_{12}$ for $z$ rotations, (iii) $J_{23}$ for $\vec{n}$ rotations, and (ii) by adjusting the ratio between them the rotation axis can be tuned within the plane of the $z$ axis and $\vec{n}$ to achieve an $x$ rotation. (b) and (c) show rotations on the Bloch sphere color coded respective to the boxes (i)-(iii).
This paper is organized as follows. We introduce the exchange-only qubit in Sec. II and theoretically show how single-qubit gates can be implemented using fewer pulses in Sec. II.1. We experimentally demonstrate the concept of simultaneous exchange pulses in Sec. II.2. We show how a simultaneous gate can perform a direct $X$ rotation by tuning the gate voltages into the appropriate regime and discuss the relative effect of charge noise in comparison to the sequential counterpart. In Sec. II.3 we investigate two-qubit gate subsequences and present an operation introducing phases between qubit states and selected leakage states. We extend our description from a linear to an all-to-all connectivity in Sec. II.4 and summarize the results for frequently used quantum gates in Tables 1 and 2.
@@ II.1 Single-qubit gates
We consider the qubit space for the decoherence-free subsystem with total spin 1/2, i.e., we do not favor one of the two possible spin polarizations. Since exchange coupling preserves the total spin, the Hamiltonian can be effectively written in the $\{|0_{\pm}\rangle,|1_{\pm}\rangle\}$ basis as
@@
In general, any rotation axis in the $x$-$z$ plane enclosed by the $z$ axis and $\vec{n}$ is possible. Thus a direct rotation around the $x$ axis in the opposite rotational direction (Fig. 1(c)) can be realized using a simultaneous (or parallel [7]) pulsing of both exchange values with $J_{23}(t)=2J_{12}(t)=2J(t)$, as the $\sigma_{z}$ terms cancel in the Hamiltonian in Eq. (3). The time evolution is then given by
@@
II.2 Experiment and simulation
Figure 2: (a) Experimental fingerpinch plot with $X$ eigenstate prepared. Barriers controlling $J_{n}$ and $J_{z}$ exchange are pulsed for a fixed time with amplitudes given by the plot axes. One minus the ground state measurement probability is plotted for different colors shown on the scale bar to the right. Simulated rotations are plotted in the upper left inlay. (b) Bloch sphere showing $J_{n}$ (orange) and $J_{x}$ (blue) rotation axes and rotation trajectories. For $J_{n}$ rotations, the solid and dashed orange circles represent the trajectory of the singlet state and $X$ eigenstate rotations, respectively. (c) Fingerpinch plot with echoing and repeated exchange pulsing during the pulse sequence. $J_{n}$ rotations (lower right region) are more visible due to the repeated pulsing and the $Z$-$\pi$ pulses rotating the eigenstate to parts of the Bloch sphere more orthogonal to $J_{n}$. (d) $J_{n}$ time-domain rotations for a fixed barrier pulse (red star in (a)) for singlet state (blue points) and $X$ eigenstate (red points). The quality factor of the rotations is 19.9. (e) $J_{x}$ time-domain rotations with singlet state (blue points) and $X$ eigenstate (red points) at a fixed barrier pulse (red triangle in (a)). The quality factor $Q=8.2$ is lower in this case. The $X$ eigenstate does not rotate, which is expected (red points).
@@ II.2 Experiment and simulation
To verify and characterize the operation of simultaneous exchange, a SiGe quantum dot device fabricated by Intel [15] is tuned up as an exchange-only qubit [16]. The qubit is encoded across three quantum dots underneath metal “plunger” gates with “barrier” gates in between dots to control the exchange. The qubit is controlled via baseband pulses on the barrier gates and readout is performed using Pauli spin blockade with signal amplified by SiGe heterojunction bipolar transistor (HBT) cryogenic amplifiers [17].
@@
To reduce sensitivity to charge noise and calibrate exchange, “fingerprint” plots [18] are acquired for the “$J_{z}$” and “$J_{n}$” control axes, corresponding to $J_{12}$ and $J_{23}$, respectively, where $J_{n}$ points down 120 degrees from the $z$-axis, analogous to Fig. 1(b). For that the exchange interaction is varied as a function of barrier voltage and detuning, leading to the typical appearance of a fingerprint. In fact, for a fixed time of 100 ns, the barrier gate bias is pulsed to larger values while the voltage detuning between two surrounding plunger gates is pulsed from negative to positive values. $J_{n}$ fingerprint plots can be taken after preparing a singlet state, but $J_{z}$ fingerprints must have a pre- and post-$J_{n}$ rotation applied which brings the state to and from the $x-y$ plane of the Bloch sphere, respectively.
@@
A typical fingerprint plot uses two plungers and one barrier to increase only one exchange term and calibrate one rotation axis, which can then be used for serial exchange control. To increase both exchange terms simultaneously, the two barrier gates are pulsed together for a fixed time, and $1-P_{|0\rangle}$ is plotted in the “fingerpinch” plot in Figure 2(a), where $P_{|0\rangle}$ is the probability of measuring the ground state $|0_{\pm}\rangle$. The Pauli spin blockade readout only detects whether the two neighboring spins 1 and 2 in Fig. 1(a) are in a singlet state or not, entirely neglecting the spin projection $S_{z}$. If a singlet state is prepared, it will only be rotated by the $J_{n}$ exchange, therefore, we prepare an $X$ eigenstate using a $Y$-$\pi/2$ prerotation, which uses alternating $J_{n}$ and $J_{z}$ exchange pulses. The $X$ eigenstate can be rotated by either $J_{n}$ (lower right region of plot) or $J_{z}$ (upper left region). Additionally, the $X$ eigenstate will not be rotated when the total exchange points along the $x$-axis, therefore the darker, linear region of the upper right defines the “$J_{x}$” regime and is indicated with a white dashed line. Theoretical calculations of the expected rotation outcomes are plotted in the upper left and agree well with the data.
@@
A modified fingerpinch plot is acquired and plotted in Figure 2(c), where alternating $Z$-$\pi$ gates and exchange pulses are applied to echo the state and extend the rotation visibility, particularly in the $J_{n}$ regime (lower right region of plot). Here, the $J_{x}$ axis highlighted by the white dashed line is more visible throughout the plot with theoretical predictions again in agreement. Figure 2(d) shows time-domain data for $J_{n}$ rotations, where the barrier is pulsed to a fixed value shown by the red star in figure 2(a). Either a singlet state (blue data points) or an $X$ eigenstate (red points) is prepared as depicted as a blue and red initial state on the Bloch sphere in Fig. 2(b). The frequency of the rotations is $f=33.9$ MHz with a quality factor of $Q=19.9$. Similarly, Figure 2(e) shows the $J_{x}$ time-domain data, taken with barriers pulsed to the upper right region of the fingerpinch plot (red triangle point in Figure 2(a)). The frequency is $f=9.5$ MHz and the quality factor is a lower value $Q=8.2$. As expected, $J_{x}$ does not rotate an $X$ eigenstate (red points). The rotations corresponding to $J_{n}$ and $J_{x}$ are shown in Fig. 2(b) in orange and light blue, respectively.
@@
Charge noise enters in the Hamiltonian (1) as fluctuations in the exchange interaction. A simple estimation for the fidelity is to assume quasi-static exchange fluctuations on each exchange value separately. In Fig. 3 we compare the sequential (solid lines) and simultaneous (dashed lines) $X$ gate with linear connectivity for various $\sigma_{J_{12}}/J_{12}$, where $\sigma_{J_{12}}$ is the standard deviation. We find that both, the sequential and the simultaneous pulse gate fidelities have a similar behaviour with a stronger dependence on the fluctuations in $J_{23}$ than in $J_{12}$. For the sequential gate this can be explained by the larger overall rotation angle around $\vec{n}$ induced by $J_{23}$ in the $X$ gate construction, whereas for the simultaneous gate it is due to the fact that $J_{23}=2J_{12}$. Any deviation from this ratio results in a tilt of the rotation axis and can result in an additional over- or underrotation. For larger $\sigma_{J_{23}}/J_{23}$ we find a slightly better performance for the sequential gate, however, for biased noise in the different exchange interactions with, e.g., $\sigma_{J_{12}}/J_{12}=10$% and $\sigma_{J_{23}}/J_{23}\leq 2$%, we even find a fidelity improvement when using the simultaneous pulses.
@@
Figure 3: Calculated $X$ gate fidelity suffering from quasi-static exchange fluctuations with standard deviations $\sigma_{J_{12}}$ and $\sigma_{J_{23}}$. We compare the sequential and simultaneous $X$ gate (with linear connectivity) for various combinations. Both, the sequential and the simultaneous pulse gate fidelity strongly depends on the fluctuations in $J_{23}$, whereas the fluctuations in $J_{12}$ are rather negligible.
II.3 Two-qubit gate subsequences
@@
To construct two-qubit gates for qubits denoted $A$ and $B$, it is advantageous to use the total spin basis of the entire six spin system $|S_{\rm tot}S_{z,\rm tot}S_{A}S_{B}S_{A,1,2}S_{B,1,2}\rangle$. Since $S_{\rm tot}$ and $S_{z,\rm tot}$ are preserved under exchange interactions, blocks with different total spins are separable as they do not couple to each other [7, 10]. As shown in Fig. 6(a) we label the chain of six spins from left to right with numbers and assume the outer two spins of each triple to be initialized in a singlet configuration as described in Ref. [14]. We label the left three spins as qubit $A$ and the right three spins as qubit $B$. For this configuration we list the possible basis states in the total spin basis of the $S_{\rm tot}=0$ and 1 subspace in Fig. 6(b) and (e), respectively. Each qubit, $A$ and $B$, is operated in the single-qubit spin $S_{A}=S_{B}=1/2$ subspace, thus only the total spin subspaces $S_{\rm tot}=0$ and $S_{\rm tot}=1$ cover the computational space. For each block the Hamiltonian is schematically shown in Fig. 6(c). The computational subspace is colored in light blue and the leakage states in gray. The Hamiltonian in Eq. (1) can be written in the $S_{\rm tot}=0,1$ blocks as shown in Fig. 6(d) and 6(f), respectively.
@@
with $\phi_{B}=\frac{J_{12}+J_{23}}{2}\tau_{B}$ and $\tau_{B}=\frac{2\pi}{\sqrt{J_{45}^{2}-J_{45}J_{56}+J_{56}^{2}}}$. We note that as long as $J_{34}=0$ the Hamiltonians of qubit $A$ and $B$ commute and thus both operations can be pulsed simultaneously, if necessary. From the expressions for $\phi_{A}$ and $\phi_{B}$ we can obtain the possible correction angles in a single drive which lie between 00 and $\pi$ as depicted in the upper plot in Fig. 4(a) for a varying ratio of $J_{23}/J_{12}$. The lower plot shows the corresponding pulse time. Consequently, we find a two-step simultaneous exchange pulse sequence to replace a three-step single pulse sequence, in case the correction phase lies in the respective phase range. As an example in Fig. 4(c) the $\pi/2-\pi/2-\pi/2$ sequence is shown, which can be replaced by two simultaneous pulses. The total pulse time $\tau^{\rm sim}_{\rm tot}$ for the simultaneous case exceeds the total pulse time $\tau^{\rm seq}_{\rm tot}$ of the sequential counterpart, $2.887\pi/J_{\rm max}>1.5\pi/J_{\rm max}$. Nevertheless, due to finite fall times of the barrier voltages, an idle time $\tau_{\rm idle}$ is usually required. For the sequential gate this yields a total gate time $\tau^{\rm seq}_{\rm tot}+3\tau_{\rm idle}$ while for the simultaneous pulses the total gate time is $\tau^{\rm sim}_{\rm tot}+2\tau_{\rm idle}$. Ultimately, for a long idle time $\tau_{\rm idle}>1.387\pi/J_{\rm max}$, the simultaneous pulsing becomes beneficial compared to the sequential state of the art. Furthermore, by taking into account the correction phase in a second pulse, the three-pulse sequences in Tab. 2 of any neighboring exchange pair can be replaced. However, when replacing the three-step sequences by the two-step simultaneous pulses in, e.g., the Fong-Wandzura CZ gate sequence, we do not find an overall time improvement, since the circuit was designed and optimized for the brick structure, where only commuting exchange interactions are allowed to be driven simultaneously. Instead, our result shows that there is room for improvement of the currently shortest sequences, an insight that may trigger the exploration of new two-qubit pulse sequences optimized for simultaneous pulsing.
II.4 All-to-all connectivity

### v2

@@ Abstract
The benefit of exchange-only qubits compared to other spin qubit types is the universal control using only voltage controlled exchange interactions between neighboring spins. As a compromise, qubit operations have to be constructed from non-orthogonal rotation axes of the Bloch sphere and result in rather long pulsing sequences. This paper aims to develop a faster implementation of single-qubit gates using simultaneous exchange pulses and manifests their potential for the construction of two-qubit gates. We introduce pulse sequences in which single-qubit gates could be executed faster and show that subsequences on three spins in two-qubit gates could be implemented in fewer steps. Our findings can particularly speed up gate sequences for realistic idle times between sequential pulses and we show that this advantage increases with more interconnectivity of the quantum dots. We further demonstrate how a phase operation can introduce a relative phase between the computational and some of the leakage states, which can be advantageous for the construction of two-qubit gates. In addition to our theoretical analysis, we experimentally demonstrate and characterize a simultaneous exchange implementation of $X$ rotations in a SiGe quantum dot device and compare to the state of the art with sequential exchange pulses.
@@ I Introduction
In the EO architecture, each qubit is represented by the spins of three electrons hosted in three adjacent quantum dots (Fig. 1(a)), and quantum gates are generated by pulsing exchange interactions between electrons. Out of the resulting eight-dimensional Hilbert space, four states are computational states and four are leakage states. Single-qubit gates involve exchange interactions between the electrons belonging to the same qubit and are usually composed of up to four rotations around two different axes [7, 9] (Fig. 1(b)). Two-qubit gates require exchange interactions between electrons from different qubits. Non-trivial pulse sequences of several exchange pulses populating and depopulating leakage states result in gates such as the CNOT, CZ or SWAP gate [7, 10, 11, 12, 13]. With few exceptions, the sequences that are found to result in two-qubit gates, however, assume only commuting exchange pulses to be pulsed at a time. Here we investigate the simultaneous application of non-commuting exchange-pulses which will also become relevant for other exchange-based qubit implementations [14, 15, 16]. We obtain quantum gates with fewer pulses, and hence faster executions of gate sequences. We derive conditions on the exchange signals and compare the results to sequential gate operations.
Figure 1: Schematic of an exchange-only qubit encoded in three linearly connected quantum dots in the (1,1,1) charging regime formed by potentials of plunger gates (${\rm P}_{i}$ of dot $i$) and barrier gates (${\rm X}_{ij}$ between dots $i$ and $j$). (a) Qubits are initialized in the singlet state of spins 1 and 2. Single qubit rotations are realized by switching on the exchange interaction (i) $J_{12}$ for $z$ rotations, (iii) $J_{23}$ for $\vec{n}$ rotations, and (ii) by adjusting the ratio between them the rotation axis can be tuned within the circular sector enclosed by the $z$ axis and $\vec{n}$ to achieve an $x$ rotation. (b) and (c) show rotations on the Bloch sphere color coded respective to the boxes (i)-(iii).
This paper is organized as follows. We introduce the exchange-only qubit in Sec. II and theoretically show how single-qubit gates can be implemented using fewer pulses and we experimentally demonstrate the concept of simultaneous exchange pulses in Sec. II.1. We show how a simultaneous gate can perform a direct $X$ rotation by tuning the gate voltages into the appropriate regime and discuss the relative effect of charge noise in comparison to the sequential counterpart. In Sec. II.2 we investigate two-qubit gate subsequences and present an operation introducing phases between qubit states and selected leakage states. We extend our description from a linear to an all-to-all connectivity in Sec. II.3 and summarize the results for frequently used quantum gates in Tables 1 and 2.
@@ II.1 Single-qubit gates
Figure 2: (a) Experimental fingerpinch plot with $X$ eigenstate prepared. Barriers controlling $J_{n}$ ($J_{23}$) and $J_{z}$ ($J_{12}$) exchange are pulsed for a fixed time with amplitudes given by the plot axes. One minus the ground state measurement probability ($1-P_{|0\rangle}$) is plotted for different colors shown on the scale bar to the right. Simulated rotations are plotted in the upper left inlay. (b) Bloch sphere showing $J_{n}$ (orange) and $J_{x}$ (blue) rotation axes and rotation trajectories. For $J_{n}$ rotations, the solid and dashed orange circles represent the trajectory of the singlet state and $X$ eigenstate rotations, respectively. (c) Fingerpinch plot with echoing and repeated exchange pulsing during the pulse sequence. $J_{n}$ rotations (lower right region) are more visible due to the repeated pulsing and the $Z_{\pi}$ pulses rotating the eigenstate to parts of the Bloch sphere more orthogonal to $J_{n}$. (d) $J_{n}$ time-domain rotations for a fixed barrier pulse (red star in (a)) for singlet state (blue points) and $X$ eigenstate (red points). The quality factor of the rotations is 19.9. (e) $J_{x}$ time-domain rotations with singlet state (blue points) and $X$ eigenstate (red points) at a fixed barrier pulse (red triangle in (a)). The quality factor $Q=8.2$ is lower in this case. The $X$ eigenstate does not rotate, which is expected (red points).
@@
We consider the qubit space for the decoherence-free subsystem with total spin 1/2, i.e., we do not favor one of the two possible spin polarizations. Since exchange coupling preserves the total spin, the Hamiltonian can effectively be written in the $\{|0_{\pm}\rangle,|1_{\pm}\rangle\}$ basis as
@@
In general, any rotation axis in the $xz$-plane enclosed by the $z$ axis and $\vec{n}$ is possible. Thus a direct rotation around the $x$ axis in the opposite rotational direction (Fig. 1(c)) can be realized using a simultaneous (or parallel [7]) pulsing of both exchange values with $J_{23}(t)=2J_{12}(t)=2J(t)$, as the $\sigma_{z}$ terms cancel in the Hamiltonian in Eq. (3). The time evolution is then given by
@@
To verify and characterize the operation of simultaneous exchange, a SiGe quantum dot device fabricated by Intel [18] is tuned up as an exchange-only qubit [19]. The qubit is encoded across three quantum dots underneath metal “plunger” gates with “barrier” gates in between dots to control the exchange. The qubit is controlled via baseband pulses on the barrier gates and readout is performed using Pauli spin blockade with signal amplified by SiGe heterojunction bipolar transistor (HBT) cryogenic amplifiers [20, 21].
@@
To reduce sensitivity to charge noise and calibrate exchange, “fingerprint” plots [22] are acquired for the “$J_{z}$” and “$J_{n}$” control axes, corresponding to $J_{12}$ and $J_{23}$, respectively, where $J_{n}$ points down 120 degrees from the $z$-axis, analogous to Fig. 1(b). For that the exchange interaction is varied as a function of barrier voltage and detuning, leading to the typical appearance of a fingerprint. In fact, for a fixed time of 100 ns, the barrier gate bias is pulsed to larger values while the voltage detuning between two surrounding plunger gates is pulsed from negative to positive values. $J_{n}$ fingerprint plots can be taken after preparing a singlet state, but $J_{z}$ fingerprints must have a pre- and post-$J_{n}$ rotation applied which brings the state to and from the $xy$-plane of the Bloch sphere, respectively.
@@
A typical fingerprint plot uses two plungers and one barrier to increase only one exchange term and calibrate one rotation axis, which can then be used for serial exchange control. To increase both exchange terms simultaneously, the two barrier gates are pulsed together for a fixed time, and $1-P_{|0\rangle}$ is plotted in the “fingerpinch” plot in Figure 2(a), where $P_{|0\rangle}$ is the probability of measuring the ground state $|0_{\pm}\rangle$. The Pauli spin blockade readout only detects whether the two neighboring spins 1 and 2 in Fig. 1(a) are in a singlet state or not, entirely neglecting the spin projection $S_{z}$. If a singlet state is prepared, it will only be rotated by the $J_{n}$ exchange, therefore, we prepare an $X$ eigenstate using a $Y_{\pi/2}$ prerotation, which uses alternating $J_{n}$ and $J_{z}$ exchange pulses [9]. The $Y_{\pi/2}$ pulse is composed of one $J_{n}$ pulse that rotates the singlet state to the equator of the Bloch sphere followed by a $J_{z}$ pulse which rotates the state to the $X$ eigenaxis. The $X$ eigenstate can be rotated by either $J_{n}$ (lower right region of plot) or $J_{z}$ (upper left region). Additionally, the $X$ eigenstate will not be rotated when the total exchange points along the $x$-axis, therefore the darker, linear region of the upper right defines the “$J_{x}$” regime and is indicated with a white dashed line. Theoretical calculations of the expected rotation outcomes are plotted in the upper left and agree well with the data.
@@
A modified fingerpinch plot is acquired and plotted in Figure 2(c), where alternating $Z_{\pi}$ gates and exchange pulses are applied to echo the state and extend the rotation visibility, particularly in the $J_{n}$ regime (lower right region of plot). Here, the $J_{x}$ axis highlighted by the white dashed line is more visible throughout the plot with theoretical predictions again in agreement. Figure 2(d) shows time-domain data for $J_{n}$ rotations, where the barrier is pulsed to a fixed value shown by the red star in figure 2(a). Either a singlet state (blue data points) or an $X$ eigenstate (red points) is prepared as depicted as a blue and red initial state on the Bloch sphere in Fig. 2(b). The frequency of the rotations is $f=33.9$ MHz with a quality factor of $Q=19.9$. The quality factor, $Q$, is defined as the characteristic decay time of an exponentially decaying sinusoid fit to the rotation data multiplied by the frequency of the rotations. Similarly, Figure 2(e) shows the $J_{x}$ time-domain data, taken with barriers pulsed to the upper right region of the fingerpinch plot (red triangle point in Figure 2(a)). The frequency is $f=9.5$ MHz and the quality factor is a lower value $Q=8.2$. As expected, $J_{x}$ does not rotate an $X$ eigenstate (red points). The rotations corresponding to $J_{n}$ and $J_{x}$ are shown in Fig. 2(b) in orange and light blue, respectively.
@@
Charge noise enters in the Hamiltonian (1) as fluctuations in the exchange interaction. A simple estimation for the fidelity is to assume quasi-static exchange fluctuations on each exchange value separately, i.e. $J_{ij}\rightarrow J_{ij}+\delta J_{ij}$, where $\delta J_{ij}$ is a Gaussian-distributed exchange fluctuation with standard deviation $\sigma_{J_{ij}}$. In Fig. 3 we compare the sequential (solid lines) and simultaneous (dashed lines) $X$ gate with linear connectivity for various $\sigma_{J_{12}}/J_{12}$, where $\sigma_{J_{12}}$ is the standard deviation for fluctuations in $J_{12}$. We find that both, the sequential and the simultaneous pulse gate fidelities have a similar behaviour with a stronger dependence on the fluctuations in $J_{23}$ than in $J_{12}$. For the sequential gate this can be explained by the larger overall rotation angle around $\vec{n}$ induced by $J_{23}$ in the $X$ gate construction, whereas for the simultaneous gate it is due to the fact that $J_{23}=2J_{12}$. Any deviation from this ratio results in a tilt of the rotation axis and can result in an additional over- or underrotation. For larger $\sigma_{J_{23}}/J_{23}$ we find a slightly better performance for the sequential gate, however, for biased noise in the different exchange interactions with, e.g., $\sigma_{J_{12}}/J_{12}=10$% and $\sigma_{J_{23}}/J_{23}\leq 2$%, we even find a fidelity improvement when using the simultaneous pulses. Typical exchange fluctuations $\sigma_{J_{ij}}/J_{ij}$ in the experiment in Fig. 2 correspond to 0.01 to 0.03.
@@
Figure 3: Calculated $X$ gate fidelity suffering from quasi-static exchange fluctuations with standard deviations $\sigma_{J_{12}}$ and $\sigma_{J_{23}}$. We compare the sequential and simultaneous $X$ gate (with linear connectivity) for various combinations. Both, the sequential and the simultaneous pulse gate fidelity strongly depend on the fluctuations in $J_{23}$, whereas the fluctuations in $J_{12}$ are rather negligible.
II.2 Two-qubit gate subsequences
@@
To construct two-qubit gates for two qubits $A$ and $B$, it is advantageous to use the total spin basis of the entire six spin system $|S_{\rm tot}S_{z,\rm tot}S_{A}S_{B}S_{A,1,2}S_{B,1,2}\rangle$. Since $S_{\rm tot}$ and $S_{z,\rm tot}$ are preserved under exchange interactions, blocks with different total spins are separable as they do not couple to each other [7, 10]. As shown in Fig. 6(a) we label the chain of six spins from left to right with numbers and assume the outer two spins of each triple to be initialized in a singlet configuration as described in Ref. [17]. We label the left three spins as qubit $A$ and the right three spins as qubit $B$. For this configuration we list the possible basis states in the total spin basis of the $S_{\rm tot}=0$ and 1 subspace in Fig. 6(b) and (e), respectively. Each qubit, $A$ and $B$, is operated in the single-qubit spin $S_{A}=S_{B}=1/2$ subspace, thus only the total spin subspaces $S_{\rm tot}=0$ and $S_{\rm tot}=1$ cover the computational space. For each block the Hamiltonian is schematically shown in Fig. 6(c). The computational subspace is colored in light blue and the leakage states in gray. The Hamiltonian in Eq. (1) can be written in the $S_{\rm tot}=0,1$ blocks as shown in Fig. 6(d) and 6(f), respectively.
@@
with $\phi_{B}=\frac{J_{12}+J_{23}}{2}\tau_{B}$ and $\tau_{B}=\frac{2\pi}{\sqrt{J_{45}^{2}-J_{45}J_{56}+J_{56}^{2}}}$. We note that as long as $J_{34}=0$ the Hamiltonians of qubit $A$ and $B$ commute and thus both operations can be pulsed simultaneously, if necessary. From the expressions for $\phi_{A}$ and $\phi_{B}$ we can obtain the possible correction angles in a single drive which lie between 00 and $\pi$ as depicted in the upper plot in Fig. 4(a) for a varying ratio of $J_{23}/J_{12}$. The lower plot shows the corresponding pulse time. Consequently, we find a two-step simultaneous exchange pulse sequence to replace a three-step single pulse sequence, in case the correction phase lies in the respective phase range. As an example in Fig. 4(c) the $\pi/2-\pi/2-\pi/2$ sequence is shown, which can be replaced by two simultaneous pulses. The total pulse time $\tau^{\rm sim}_{\rm tot}$ for the simultaneous case exceeds the total pulse time $\tau^{\rm seq}_{\rm tot}$ of the sequential counterpart, $2.887\pi/J_{\rm max}>1.5\pi/J_{\rm max}$. Nevertheless, due to finite fall times of the barrier voltages, an idle time $\tau_{\rm idle}$ is usually required. For the sequential gate this yields a total gate time $\tau^{\rm seq}_{\rm tot}+3\tau_{\rm idle}$ while for the simultaneous pulses the total gate time is $\tau^{\rm sim}_{\rm tot}+2\tau_{\rm idle}$. Ultimately, for a long idle time $\tau_{\rm idle}>1.387\pi/J_{\rm max}$, the simultaneous pulsing becomes beneficial compared to the sequential state of the art. Furthermore, by taking into account the correction phase in a second pulse, the three-pulse sequences in Tab. 2 of any neighboring exchange pair can be replaced. In Appendix B we exemplarily replace the three-step sequences with the two-step simultaneous pulses in the Fong-Wandzura CZ gate sequence, calculate the gate times, and estimate the gate performance. However, there we do not find an overall time improvement, since the circuit was designed and optimized for the brick structure, where only commuting exchange interactions are allowed to be driven simultaneously. Instead, our result shows that there is room for improvement of the currently shortest sequences, an insight that may trigger the exploration of new two-qubit pulse sequences optimized for simultaneous pulsing.
II.3 All-to-all connectivity